OpenStudy (mellamokatie):

5. Multiply 7z^3(2z^2-4z). A.14z^5-28z^4 B. 9z^5-11z&4 C. 14z^6-28z^4 D. -14z

OpenStudy (welshfella):

do you know how to multiply exponential terms? eg a^2 * a^3 = a^(2+3) = a^5

OpenStudy (mellamokatie):

Yes

OpenStudy (welshfella):

Ok you need also to know the distributive property a(b + c) = ab + ac apply these 2 rules and you have the answer (OH - you need to know how to multiply numbers as well!)

OpenStudy (mellamokatie):

Oh okay thank you give me one second!

OpenStudy (mellamokatie):

@welshfella I would add 3 and 2 right?

OpenStudy (mellamokatie):

@DopeBoyDee

OpenStudy (mellamokatie):

I've gotten \[14z^5 \]

OpenStudy (mellamokatie):

But that's not the final answer

OpenStudy (mellamokatie):

Wait its A

OpenStudy (welshfella):

14z^5 is correct for the first term

OpenStudy (welshfella):

and second term = 7z^3 * -4z

OpenStudy (mellamokatie):

3z^3?

OpenStudy (welshfella):

= -28z^4

OpenStudy (mellamokatie):

OHHH wait multiply right

OpenStudy (welshfella):

yes

OpenStudy (mellamokatie):

Thank you!

OpenStudy (welshfella):

-Yes - dont forget the Distriuutive Property

OpenStudy (mellamokatie):

Right

OpenStudy (welshfella):